40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Apr 27 19:37:49 2021, shruti, Update, PSOMA, PDH error signal PDHerrorsig.mp4BeatNX20210427.pdfIMG_0215.jpgPDH.pdf
    Reply  Tue Apr 27 20:24:18 2021, rana, Update, PSOMA, PDH error signal 
    Reply  Fri May 7 13:15:07 2021, shruti, Update, PSOMA, Cavity locking, feedback to laser IMG_0217.pdf
       Reply  Mon May 10 10:20:54 2021, shruti, Update, PSOMA, Cavity locking, feedback to laser ControlSigLockedCav.pdfControlSigCavLocked2.pdf
          Reply  Mon May 10 15:41:41 2021, rana, Update, PSOMA, Cavity locking, feedback to laser 
Message ID: 2731     Entry time: Fri May 7 13:15:07 2021     In reply to: 2713     Reply to this: 2732
Author: shruti 
Type: Update 
Category: PSOMA 
Subject: Cavity locking, feedback to laser 

[Shruti, Rana]

I made that the ITC 502 current driver was in constant current mode where the feedback coeff is 20 mA / V \pm 5% and bandwidth 500 kHz; with the laser coefficient being 100 MHz the net feedback coefficient would be 2 GHz/V.Then, I looked at the level of the PDH signal which was roughly \pm 1V.

In order to attenuate the signal going in to the MOD IN of the ITC 502 current driver, Rana suggested using the 50 Ohm output of the SR560 with two series 20 dB attenuators and a BNC T with one port going to the MOD IN and the other port terminated with 50 Ohm.

With the SR 560 set with 1 Hz low pass and gain of 10, the cavity locked for 3-10 seconds when it came close to resonance as seen on the monitor.

The three attachments are the PDH signal (purple), cavity reflection (blue), and cavity transmission (yellow) when the cavity seemed to lock.

Attachment 1: IMG_0217.pdf  22.935 MB  Uploaded Fri May 7 14:15:36 2021  | Hide | Hide all
IMG_0217.pdf
ELOG V3.1.3-